# High-energy physics image classification: A survey of Jet applications

### v1

@@ 5 HEP Jet image classification
ML, especially neural networks, has a rich historical presence in the field of particle physics. The concept of applying neural networks for tasks like distinguishing quarks and gluons, tagging Higgs particles, and identifying particle tracks has been around for more than two and a half decades. Nevertheless, the recent advancements in deep learning and the increased computational capabilities offered by graphics processing units (GPUs) have led to a significant enhancement in image recognition technology. As a result, there has been a renewed and heightened interest in utilizing these techniques. In the subsequent sections, we provide an overview of SOTA methods in both ML and DL. Figure 13 depicts a taxonomy of existing ML and DL techniques (discussed in Section 5), summarizes the previously discussed AI-based Jet images models (discussed in Section 4), preprocessing and datasets (discussed in Section 3), and metrics (discussed in Section 2).
@@ 5.2 DNN-based
The identification of b Jets using QCD-inspired observables was studied in [32]. It involves the application of Jet substructure observables, such as one-dimensional Jet angularities and the two-dimensional primary lund plane (PLP). DNNs are employed to identify b Jets using these QCD-inspired observables. The DNNs are trained on a set of input features, which include Jet angularities and the PLP, in order to efficiently distinguish b Jets from light ones. The performance of the DNNs is evaluated by comparing their results with those of conventional track-based taggers, such as JetFitter, IP3D, and DL1 taggers. In this study, the results indicate that the DNN discriminants exhibit better performance than the IP3D tagger. The researchers in [33] discusses the application of DNNs to a wide range of physics problems, particularly in HEP. Specifically, DNNs have been successfully applied to tasks such as Jet tagging and event classification. The authors explore the use of a simple but effective preprocessing step that transforms observational quantities into a binary number with a fixed number of digits, representing the quantity or magnitude in different scales. This approach has been shown to significantly improve the performance of DNNs for specific tasks without complicating feature engineering, particularly in b-Jet tagging using daughter particles’ momenta and vertex information.
@@
5.3 CNN-based
CNNs have revolutionized Jet image classification and prediction in particle physics. CNNs excel in image recognition by leveraging convolutional layers, weight sharing, and pooling to capture hierarchical features, enabling effective pattern recognition and classification [35, 36]. This enables precise particle identification using Jet images, improved event classification, and deeper insights into HEP experiments, advancing researchers’ understanding of fundamental particles and interactions. For example, the study [37] demonstrates CNN’s efficacy in predicting energy loss for quark and gluon Jets, yielding comparable results. It highlights distinctions post-quenching and employs deep learning for classification, emphasizing energy loss’s impact on classification difficulty. The authors in [38] investigate the capability of CNNs in discriminating quark and gluon Jets, comparing their performance to traditionally designed physics observables. Similarly, in order to discriminate Quark-Gluon Jet, Lee et al. in their research [39], employed various pretrained CNN models, including VGG, ResNet, Inception-ResNet, DenseNet, Xception, Vanilla ConvNet, and Inception-ResNet, to classify Jet images for distinguishing quark and gluon hadron Jets. The study reveals that DenseNet outperforming larger, higher-structured networks. Despite marginal improvements over a traditional BDT classifier, stability in training can be enhanced using the RMSProp optimizer, an adaptive learning rate optimization algorithm.
@@ 5.3 CNN-based
5.4 Adversarial training-based
@@ 5.6 Other
Table 6: Summary of the performance of certain deep learning frameworks proposed for HEP. Only the best performance is reported in the case of multiple tests.
Ref.
DLM
Dataset
Description
BP (%)
Limitations
PLA
[1]
CNN
QCD multi-jet
Classification of multi-Jet events using CNN at high energies of 13 TeV
AUC= 99.03
The proposed CNN model needs validation with additional datasets to ensure its generalizability.
No
[17]
SVM
Simulated
BIP features invariant under boosts for improved Jet tagging
Acc= 92.7
Performance could be enhanced through comprehensive hyperparameter tuning.
Yes [^12] https://zenodo.org/records/7271316
[25]
DNN
Simulated
Sequence of jet components arranged in a specific order for training inputs.
Acc= 50
Could be enhanced by employing the LSTM method to efficiently classify Jet from background.
No
[30]
DNN
Simulated [^13] https://www.igb.uci.edu/~pfbaldi/physics/
DNNs for categorizing Jet substructure in HEP
AUC= 95.3
The accuracy of the DNN models is limited by the accuracy of the simulation models used to generate the training data.
No
[31]
DNN
Higgs
Clarifying HEP event classification with SHAP
Acc= 66
SHAP may not comprehensively capture feature interactions or explain model behavior in all cases. It could demand substantial computational resources for large datasets or intricate models.
Yes [^14] https://github.com/rpezoa/hep_shap/
[32]
DNN
ATLAS
Detection of b Jets utilizing QCD-inspired measurements
AUC= 67
The DNN performed slightly less effectively than the JetFitter algorithm.
No
[34]
DNN
Pythia
@@
Creating images with low pixel density in particle physics for two cased $D+D$ and $D+C$.
AUC= 86.9,
@@
Slower than the non-autoregressive model LAGAN. $D+D$ performed better than $D+C$ for both Pythia and Monte Carlo images.
Yes [^15] http://mlphysics.ics.uci.edu/
[37]
CNN
Simulated
CNN for predicting quark and gluon Jets
Acc= 75.9
The higher the energy loss, the more challenging the task of classifying the Jets becomes.
No
[47]
RecNN
Simulated
Enhance Quark/gluon classification
AUC= 86.37
Event-level analysis is not performed.
Yes [^16] https://github.com/glouppe/recnn
[49]
CNN and AE
Daya Bay
Classification for different event types, including IBD prompt, IBD delay, Muon, Flasher, and other
Acc= 99.9 (Muon)
SVM and KNN exhibit inferior performance compared to CNN in identifying event types. Moreover, semi-supervised techniques have not been examined.
No
[50]
CNN
Simulated
Employing a quantum CNN to categorize events in HEP.
Acc= 97.5
Quantum CNN showed a lower performance than CNN when it comes to a binary classification of Muon and Electron. Besides, CNN showed low performance when classifying Muon and Pion compared to quantum CNN.
No
[51]
ML
ATLAS
Predict if the LHC trials have dismissed a new physics model
Acc= 93.8
Enhancing reliability can be achieved by requiring a minimum confidence level for the prediction.
Yes [^17] http://susyai.hepforge.org
[52]
ANN
Simulated
Identifying boosted top quarks using pattern recognition through an ANN in HEP experiments.
Eff= 60
It has 4% mis-tag rate. It exclusively utilizes hadronic calorimete (HCAL) data, though additional data, like sub-jet b-tags, are crucial for top tagging.
No
[53]
DNN
Real data
Enhancing jet reconstruction at CMS through deep learning
FPR= 65
The computational costs, wnen employing the proposed model, have not been verified.
No
Abbreviations: Deep learning model (DLM); Best performance (BP); Project link availability (PLA).
@@ 6 Applications of AI-based Jet image
Jet images processed through ML and DL techniques hold vast potential across various applications within the HEP domain, some of theme are already described in [2]. This section presents a comprehensive overview of cutting-edge work in this area, categorized into several key domains: Jet parameter scanning, event classification, Jet tagging, multi-Jet classification, energy estimation, and beyond. The taxonomy of AI-based jet image applications is visualized in Figure 17, illustrating their scope and relationships. The section thoroughly reviews some applications conducted by researchers, while suggesting future directions for those not yet explored. Additionally, Table 6 provides a concise summary of performance metrics, limitations, online project availability, and results obtained across these applications, offering valuable insights into their efficacy and applicability.
[54]
@@ 6.2 Jet classification and tagging
Particles generated in a collider with significant center-of-mass energy typically exhibit high velocity. As a result, their decay products tend to align closely, leading to overlapping jets. It is crucial in collider data analysis to discern whether a jet originates from a solitary light particle or from the decay of a heavier particle. Traditional approaches rely on manually crafted distribution features based on energy deposition in calorimeter cells. However, due to the intricate nature of the data, ML techniques have proven more efficient than human efforts for this task [58]. In [59], the Jet image concept treats the detector as a camera, capturing Jet energy distribution in calorimeters as a digital image. This enables Jet tagging as a pattern recognition task, utilizing machine learning, like Fisher classification, to differentiate between hadronic W boson decay and Jets from quarks or gluons. Monte Carlo simulation shows superior discrimination compared to traditional methods, offering insights into Jet structure. In [38], CNNs improve tagging by treating Jet energy distribution as an image, using channels for features like particle momentum and count. Results show CNNs can surpass traditional methods, providing reliable insights from collider simulation data despite variations in event generators. However, CNNs demonstrate a lack of sensitivity to quark/gluon Jets from different generators, akin to conventional Jet measurements. Moving on, in [60], Jet tagging is performed using RNN, leveraging the similarity between Jet clustering and natural language structure. Final-state particle four-momenta are treated as language words, and Jet clustering as grammatical analysis. RNN efficiently processes the tree-like Jet structures, enabling direct use of particle data regardless of count. This method yields higher data utilization efficiency and prediction accuracy than Jet image-based ML, extending to event classification. In [47], RNNs distinguish quark and gluon Jets, showing higher gluon suppression. Factors affecting RNN performance are explored, with preliminary quark tagging results. Numerous explorations for phenomena beyond the SM at the LHC depend on top tagging techniques that distinguish between boosted hadronic top quarks and the more prevalent jets that originate from light quarks and gluons. The HCAL essentially captures a "digital image" of each jet, where the pixel brightness represents the energy deposited in HCAL cells. Therefore, top tagging is essentially a matter of recognizing patterns. The work in [52] propose a novel top tagging algorithm based on an ANN, a popular pattern recognition approach. The ANN is developed using a substantial dataset of boosted tops along with light quark/gluon jets and is subsequently evaluated on separate datasets. In Monte Carlo simulations, particularly within the 1100-1200 GeV range, the ANN-based tagger demonstrates outstanding efficacy.
@@ 6.4 Jet image generation
In order to study new physics phenomena at the LHC, it is necessary to simulate Monte Carlo events for both new physics signals and backgrounds. This simulation helps predict the experimental data expected from collider experiments. However, generating the large number of simulated events required for data analysis is time-consuming and computationally intensive using existing algorithms. Additionally, accurately simulating how energetic particles interact with detector materials can be a time-consuming process. In [61], researchers proposed using GANs to build LAGAN framework, that is trained to generate authentic radiation distributions from simulated collisions involving high-energy particles. The authors found that the generated Jet images exhibited a wide range of pixel brightness levels and accurately reproduced low-dimensional physical observables such as reconstructed Jet mass and n-subjettiness. However, the study also acknowledges the limitations of this method and presents an empirical validation of the image quality. With further improvement, this approach could lead to faster simulation of HEPs events. Physicists at the LHC use complex simulations to predict experimental outcomes. Generating vast amounts of simulated data is costly, but crucial for technique development. Challenges include accurately modeling detectors and particle interactions. In [62], researchers proposed a GAN-nased model for fast, accurate simulation of electromagnetic calorimeters. Despite ongoing precision challenges, our solution offers significant speed-ups, up to 100,000×, promising savings in computing resources and advancing physics research at the LHC and beyond.
@@ 7 Future direction and outlook
Transfer learning (TL) detailed in [63, 64, 65], is poised to revolutionize Jet HEP applications by leveraging pre-trained models from vast datasets to enhance performance on specific tasks with limited data. This approach can significantly reduce computational costs and training times, making it ideal for adapting models to new experiments or rare phenomena. As HEP experiments generate increasingly complex data, the ability to apply knowledge from one context to another will be invaluable for improving event classification, anomaly detection, and signal processing. Looking ahead, transfer learning will be crucial for efficiently extracting insights from new particle interactions and advancing our understanding of fundamental physics. Exploring advanced CNN architectures as source of prior knowledge like AlexNet, Fast-R-CNN, VGG, ResNet, GoogLeNet could offer potential improvements to target model that conduct many AI-based Jet tasks [24]. Generalizing the top tagger to classify other boosted objects, such as W/Z bosons, Higgs bosons, and BSM particles, is straightforward and extending it to partially-merged and fully resolved tops could enhance background rejection.

### v2

@@ 5 HEP Jet image classification
ML, especially neural networks, has a rich historical presence in the field of particle physics. The concept of applying neural networks for tasks like distinguishing quarks and gluons, tagging Higgs particles, and identifying particle tracks has been around for more than two and a half decades. Nevertheless, the recent advancements in DL and the increased computational capabilities offered by graphics processing units (GPUs) have led to a significant enhancement in image recognition technology. As a result, there has been a renewed and heightened interest in utilizing these techniques. In the subsequent sections, we provide an overview of SOTA methods in both ML and DL. Figure 13 depicts a taxonomy of existing ML and DL techniques (discussed in Section 5), summarizes the previously discussed AI-based Jet images models (discussed in Section 4), preprocessing and datasets (discussed in Section 3), and metrics (discussed in Section 2).
@@ 5.2 DNN-based methods
In the study conducted by [49], the identification of $b$ Jets was investigated utilizing QCD inspired observables. The process entails the utilization of Jet substructure observables, including one-dimensional Jet angularities and the two-dimensional primary lund plane (PLP). DNNs are employed to identify $b$ Jets using these QCD-inspired observables. The DNNs are trained on a set of input features, which include Jet angularities and the PLP, in order to efficiently distinguish $b$ Jets from light ones. The performance of the DNNs is evaluated by comparing their results with those of conventional track-based taggers, such as JetFitter, IP3D, and DL1 taggers. In this study, the results indicate that the DNN discriminants exhibit better performance than the IP3D tagger. The researchers in [50] discusses the application of DNNs to a wide range of physics problems, particularly in HEP. Specifically, DNNs have been successfully applied to tasks such as Jet tagging and event classification. The authors explore the use of a simple but effective preprocessing step that transforms observational quantities into a binary number with a fixed number of digits, representing the quantity or magnitude in different scales. This approach has been shown to significantly improve the performance of DNNs for specific tasks without complicating feature engineering, particularly in b-Jet tagging using daughter particles’ momenta and vertex information.
@@
5.3 CNN-based methods
CNNs have revolutionized Jet image classification and prediction in particle physics. CNNs excel in image recognition by leveraging convolutional layers, weight sharing, and pooling to capture hierarchical features, enabling effective pattern recognition and classification [52, 53]. This enables precise particle identification using Jet images, improved event classification, and deeper insights into HEP experiments, advancing researchers’ understanding of fundamental particles and interactions. For example, the study [54] demonstrates CNN’s efficacy in predicting energy loss for quark and gluon Jets, yielding comparable results. It highlights distinctions post-quenching and employs DL for classification, emphasizing energy loss’s impact on classification difficulty. The authors in [55] investigate the capability of CNNs in discriminating quark and gluon Jets, comparing their performance to traditionally designed physics observables. Similarly, in order to discriminate Quark-Gluon Jet, Lee et al. in their research [56], employed various pretrained CNN models, including VGG, ResNet, Inception-ResNet, DenseNet, Xception, Vanilla ConvNet, and Inception-ResNet, to classify Jet images for distinguishing quark and gluon hadron Jets. The study reveals that DenseNet outperforming larger, higher-structured networks. Despite marginal improvements over a traditional BDT classifier, stability in training can be enhanced using the RMSProp optimizer, an adaptive learning rate optimization algorithm.
@@ 5.3 CNN-based methods
5.4 Adversarial training-based methods
@@ 5.6 MLP-based methods
Table 6: Summary of the performance of certain DL frameworks proposed for HEP. Only the best performance is reported in the case of multiple tests.
Ref. DLM Dataset Description BP (%) Limitations PLA
[16] CNN QCD multi-jet Classification of multi-Jet events using CNN at high energies of 13 TeV AUC= 99.03 The proposed CNN model needs validation with additional datasets to ensure its generalizability. No
[34] SVM Simulated BIP features invariant under boosts for improved Jet tagging Acc= 92.7 Performance could be enhanced through comprehensive hyperparameter tuning. Yes [^12] https://zenodo.org/records/7271316
[42] DNN Simulated Sequence of jet components arranged in a specific order for training inputs. Acc= 50 Could be enhanced by employing the LSTM method to efficiently classify Jet from background. No
[47] DNN Simulated [^13] https://www.igb.uci.edu/~pfbaldi/physics/ DNNs for categorizing Jet substructure in HEP AUC= 95.3 The accuracy of the DNN models is limited by the accuracy of the simulation models used to generate the training data. No
[48] DNN Higgs Clarifying HEP event classification with SHAP Acc= 66 SHAP may not comprehensively capture feature interactions or explain model behavior in all cases. It could demand substantial computational resources for large datasets or intricate models. Yes [^14] https://github.com/rpezoa/hep_shap/
[49] DNN ATLAS Detection of b Jets utilizing QCD-inspired measurements AUC= 67 The DNN performed slightly less effectively than the JetFitter algorithm. No
[51] DNN Pythia
@@
Creating images with low pixel density in particle physics for two cased $D+D$ and $D+C$. AUC= 86.9,
@@
Slower than the non-autoregressive model LAGAN. $D+D$ performed better than $D+C$ for both Pythia and Monte Carlo images. Yes [^15] http://mlphysics.ics.uci.edu/
[54] CNN Simulated CNN for predicting quark and gluon Jets Acc= 75.9 The higher the energy loss, the more challenging the task of classifying the Jets becomes. No
[64] RecNN Simulated Enhance Quark/gluon classification AUC= 86.37 Event-level analysis is not performed. Yes [^16] https://github.com/glouppe/recnn
[66] CNN-AE Daya Bay Classification for different event types, including IBD prompt, IBD delay, Muon, Flasher, and other Acc= 99.9 (Muon) SVM and KNN exhibit inferior performance compared to CNN in identifying event types. Moreover, semi-supervised techniques have not been examined. No
[67] CNN Simulated Employing a quantum CNN to categorize events in HEP. Acc= 97.5 Quantum CNN showed a lower performance than CNN when it comes to a binary classification of Muon and Electron. Besides, CNN showed low performance when classifying Muon and Pion compared to quantum CNN. No
[68] ML ATLAS Predict if the LHC trials have dismissed a new physics model Acc= 93.8 Enhancing reliability can be achieved by requiring a minimum confidence level for the prediction. Yes [^17] http://susyai.hepforge.org
[69] ANN Simulated Identifying boosted top quarks using pattern recognition through an ANN in HEP experiments. Eff= 60 It has 4% mis-tag rate. It exclusively utilizes hadronic calorimete (HCAL) data, though additional data, like sub-jet b-tags, are crucial for top tagging. No
[70] DNN Real data Enhancing jet reconstruction at CMS through deep learning FPR= 65 The computational costs, wnen employing the proposed model, have not been verified. No
[71] CNN Simulated Detection of jet quenching effects caused by the presence of the quark-gluon plasma (QGP). AUC= 75 The computational costs, when employing the proposed model, have not been verified. When data normalized, AUC reached only 67% (when $p_{T,jet}>30GeV$). No
Abbreviations: Deep learning model (DLM); Best performance (BP); Project link availability (PLA); auto-encoder (AE).
@@ 6 Applications of AI-based Jet image
Jet images processed through ML and DL techniques hold vast potential across various applications within the HEP domain, some of theme are already described in [17]. This section presents a comprehensive overview of cutting-edge work in this area, categorized into several key domains: Jet parameter scanning, event classification, Jet tagging, multi-Jet classification, energy estimation, and beyond [72]. The taxonomy of AI-based jet image applications is visualized in Figure 17, illustrating their scope and relationships. The section thoroughly reviews some applications conducted by researchers, while suggesting future directions for those not yet explored. Additionally, Table 6 provides a concise summary of performance metrics, limitations, online project availability, and results obtained across these applications, offering valuable insights into their efficacy and applicability.
@@ 6.2 Jet classification and tagging
Particles generated in a collider with significant center-of-mass energy typically exhibit high velocity. As a result, their decay products tend to align closely, leading to overlapping jets. It is crucial in collider data analysis to discern whether a jet originates from a solitary light particle or from the decay of a heavier particle. Traditional approaches rely on manually crafted distribution features based on energy deposition in calorimeter cells. However, due to the intricate nature of the data, ML techniques have proven more efficient than human efforts for this task [76]. In [77], the Jet image concept treats the detector as a camera, capturing Jet energy distribution in calorimeters as a digital image. This enables Jet tagging as a pattern recognition task, utilizing ML, like Fisher classification, to differentiate between hadronic W boson decay and Jets from quarks or gluons. Monte Carlo simulation shows superior discrimination compared to traditional methods, offering insights into Jet structure. In [55], CNNs improve tagging by treating Jet energy distribution as an image, using channels for features like particle momentum and count. Results show CNNs can surpass traditional methods, providing reliable insights from collider simulation data despite variations in event generators. However, CNNs demonstrate a lack of sensitivity to quark/gluon Jets from different generators, akin to conventional Jet measurements. Moving on, in [78], Jet tagging is performed using RNN, leveraging the similarity between Jet clustering and natural language structure. Final-state particle four-momenta are treated as language words, and Jet clustering as grammatical analysis. RNN efficiently processes the tree-like Jet structures, enabling direct use of particle data regardless of count. This method yields higher data utilization efficiency and prediction accuracy than Jet image-based ML, extending to event classification. In [64], RNNs distinguish quark and gluon Jets, showing higher gluon suppression. Factors affecting RNN performance are explored, with preliminary quark tagging results. Numerous explorations for phenomena beyond the SM at the LHC depend on top tagging techniques that distinguish between boosted hadronic top quarks and the more prevalent jets that originate from light quarks and gluons. The HCAL essentially captures a "digital image" of each jet, where the pixel brightness represents the energy deposited in HCAL cells. Therefore, top tagging is essentially a matter of recognizing patterns. The work in [69] propose a novel top tagging algorithm based on an ANN, a popular pattern recognition approach. The ANN is developed using a substantial dataset of boosted tops along with light quark/gluon jets and is subsequently evaluated on separate datasets. In Monte Carlo simulations, particularly within the 1100-1200 GeV range, the ANN-based tagger demonstrates outstanding efficacy.
@@ 6.4 Jet image generation
In order to study new physics phenomena at the LHC, it is necessary to simulate Monte Carlo events for both new physics signals and backgrounds. This simulation helps predict the experimental data expected from collider experiments. However, generating the large number of simulated events required for data analysis is time-consuming and computationally intensive using existing algorithms. Additionally, accurately simulating how energetic particles interact with detector materials can be a time-consuming process. In [79], researchers proposed using GANs to build LAGAN framework, that is trained to generate authentic radiation distributions from simulated collisions involving high-energy particles. The authors found that the generated Jet images exhibited a wide range of pixel brightness levels and accurately reproduced low-dimensional physical observables such as reconstructed Jet mass and n-subjettiness. However, the study also acknowledges the limitations of this method and presents an empirical validation of the image quality. With further improvement, this approach could lead to faster simulation of HEPs events. Physicists at the LHC use complex simulations to predict experimental outcomes. Generating vast amounts of simulated data is costly, but crucial for technique development. Challenges include accurately modeling detectors and particle interactions. In [80], researchers proposed a GAN-nased model for fast, accurate simulation of electromagnetic calorimeters. Despite ongoing precision challenges, this solution offers significant speed-ups, up to 100,000×, promising savings in computing resources and advancing physics research at the LHC and beyond.
@@ 7 Future direction and outlook
Transfer learning (TL) is poised to revolutionize Jet HEP applications by leveraging pre-trained models from vast datasets to enhance performance on specific tasks with limited data. This approach can significantly reduce computational costs and training times, making it ideal for adapting models to new experiments or rare phenomena. As HEP experiments generate increasingly complex data, the ability to apply knowledge from one context to another will be invaluable for improving event classification, anomaly detection, and signal processing. Looking ahead, transfer learning will be crucial for efficiently extracting insights from new particle interactions and advancing our understanding of fundamental physics. Exploring advanced CNN architectures as source of prior knowledge like AlexNet, Fast-R-CNN, VGG, ResNet, GoogLeNet could offer potential improvements to target model that conduct many AI-based Jet tasks [41]. Generalizing the top tagger to classify other boosted objects, such as W/Z bosons, Higgs bosons, and BSM particles, is straightforward and extending it to partially-merged and fully resolved tops could enhance background rejection.
@@
The integration of large language models (LLMs) and generative AI into HEP has the potential to enhance the precision of particle detection and characterization. By leveraging these advanced AI models, researchers can identify subtle patterns and anomalies in jet images that might be missed by conventional methods. This improved accuracy is crucial for discovering new particles or interactions that could lead to breakthroughs in our understanding of the universe. For example, in the search for dark matter or other exotic particles, detecting faint signals amid a noisy background is a significant challenge. Generative AI can help by producing simulations that highlight these weak signals, allowing physicists to fine-tune their detection algorithms. Similarly, LLMs can assist by providing context and insight into these findings, suggesting potential theoretical implications and further areas of exploration. The application of LLMs and generative AI in HEP also promotes a more collaborative and interdisciplinary approach to research. By integrating AI experts with physicists, new methodologies and tools can be developed that leverage the strengths of both fields. This collaboration can lead to the creation of more sophisticated models that are specifically tailored to the needs of HEP. Furthermore, the insights gained from HEP research using AI can be applied to other fields, such as astrophysics, medical imaging, and materials science. This cross-pollination of ideas and techniques can drive innovation across multiple disciplines, leading to advancements that benefit a wide range of scientific endeavors.